Find n arthimrtic inserted in any two real number a&b

Grade:12

1 Answers

Arun
6 years ago
We have to insert n A.M. between a &b.
Hence total nos. of terms will be (n+2)
Now 
from Tn = a+ (n-1)*d
b = a +(n+2-1)*d
b- a =( n+1)*d
d = (b-a)/(n+1)
Hence now we can find all the mean
1st A.M = a +d ...do it yourself
2nd A.M. = a +2*d (do it yourself)